OpenStudy (abbycross167):

can someone please help me with this math question? please I'll fan and give a medal...

OpenStudy (abbycross167):

show the steps to graph this equation y > 3x - 8

OpenStudy (abbycross167):

@jim_thompson5910 can you please try to help me?

jimthompson5910 (jim_thompson5910):

first step is to replace the > with = so y > 3x - 8 becomes y = 3x - 8

jimthompson5910 (jim_thompson5910):

are you able to graph y = 3x - 8 @abbycross167 ?

OpenStudy (abbycross167):

no sir/ma'am

jimthompson5910 (jim_thompson5910):

if x is equal to say x = 0, then what must y be? replace every copy of x with 0 and evaluate to find out

jimthompson5910 (jim_thompson5910):

y = 3x-8 y = 3*0 - 8 ... replace every x with 0 y = ???

OpenStudy (abbycross167):

-8?

jimthompson5910 (jim_thompson5910):

yes so (0,-8) is one ordered pair. It's one point on the graph to find another, plug in x = 1 and determine what y must be

OpenStudy (abbycross167):

y = 3x-8 y = 3*1 - 8 ... replace every x with 1 y = 3?

jimthompson5910 (jim_thompson5910):

3*1 turns into 3 3-8 is NOT equal to 3

OpenStudy (abbycross167):

so y=5?

jimthompson5910 (jim_thompson5910):

you subtracted in the wrong order 3-8 = -5 small minus big = negative number

jimthompson5910 (jim_thompson5910):

plugging in x = 1 leads to y = -5

jimthompson5910 (jim_thompson5910):

(0,-8) is one point on the graph (1,-5) is another point on the graph plot the two points, then draw a straight line through them. Make sure the line extends as far as possible in both directions

OpenStudy (abbycross167):

okay, thank you so much I really appreciate your help!!

jimthompson5910 (jim_thompson5910):

after you get the line graphed, you'll just shade above the dotted line to graph y > 3x-8

OpenStudy (abbycross167):

yes sir thank you again!

jimthompson5910 (jim_thompson5910):

no problem